OpenStudy (anonymous):

sin^2(2 theta)=1 on the interval 0< theta< 2pi

OpenStudy (anonymous):

Well if we set \[\large 2\theta\]=a We get sin^2(a)=1 We subtract 1 from both sides to get \[\Large \sin^2(a)-1=0\] Can you factor this for me?

OpenStudy (anonymous):

\[\sin ^{2}(a)=1 \] \[\sqrt{\sin ^{2}}(a)=\sqrt{1}\] \[\sin (a)=1\] \[\sin\frac{ \pi }{ 2 }= 1\] im not to sure if this is right

OpenStudy (anonymous):

You have to factor it out since there's a sqrt

OpenStudy (anonymous):

\[\sin^2(a)-1=0~~~~~~~~~~~~~~(\sin(a)+1)*(\sin(a)-1)=0\]

OpenStudy (anonymous):

set those those two equations eqaul to 0